OpenStudy (anonymous):

solve for x cos^2 x - sin^2 x = sinx

OpenStudy (anonymous):

\[-\pi < x \le \pi\]

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

should i make everything sin?

OpenStudy (anonymous):

ok. i see where you are going with this.

OpenStudy (anonymous):

or nope don't just let me think

OpenStudy (anonymous):

then we would subtract 1 to get \[-2\sin ^{2}x = \sin x -1\] correct?

OpenStudy (anonymous):

yup it's correct but i just solved it by making everything cos

OpenStudy (anonymous):

is it easier that way?

OpenStudy (anonymous):

idk i will show u my solution

OpenStudy (anonymous):